OpenStudy (chaise):

Find the sum of the following series: 1/1+1/2+2/4+3/8+5/16+8/32+... where each numerator is a term of the Fibonacci Sequence and the denominators are in a geometrical progression.

OpenStudy (chaise):

I understand that tn=((tn-1)+(tn-2))/2

OpenStudy (chaise):

You mean (n-1)^2?

OpenStudy (anonymous):

denominator is not n^2.....it is 2^( n-1)

OpenStudy (klimenkov):

\(\frac11+\frac12+\frac24+\frac38+\frac5{16}+\frac8{32}+\cdots\)

OpenStudy (anonymous):

no..its a converging series

OpenStudy (experimentx):

sorry ..

OpenStudy (chaise):

oops, converges at 4. my bad. :s

OpenStudy (anonymous):

i think phi must be used to solve this problem

OpenStudy (klimenkov):

\(t_n=\frac {F_n}{2^n}\), where \(F_n\) - n-th term of Fibonacci sequence

OpenStudy (anonymous):

@chaise...denominator: 2^(n-1)

OpenStudy (klimenkov):

n=0,1,2,...

OpenStudy (anonymous):

Such confusion! People it's not that hard... \(\large F_n = F_{n-1} + F_{n-2}\), where n = discrete interger numbers from 1 to \(\infty\) The n-1 mean "the previous term" The n-2 means "the one before the previous term" This is recursion in action here

OpenStudy (chaise):

Right, so how do you determine the sum to infinity?

OpenStudy (anonymous):

In this case we need to modify that series so it matches the output above

OpenStudy (anonymous):

You need to do a test of some kind @chaise Alternating series test? no. P-series? no. Geometric series? no. Integral test? no. Ratio test? ;-) hmm...

OpenStudy (experimentx):

http://mathworld.wolfram.com/FibonacciNumber.html chek out (15)

mathslover (mathslover):

is it a harmonic progression ?

mathslover (mathslover):

i do think like that : this is in a harmonic progression ..

OpenStudy (experimentx):

put the value of x=1/2 divide the left by x again ... i guess this would give the answer.

OpenStudy (anonymous):

@satya_Balli I would suspect it is NOT a convergent series. Even if it's a positive number and the positive numbers go to zero they keep increasing the total sum little by little, just as the area under the following would be infinite: \[\large \int\limits_{1}^{\infty} \log |x| \ \ dx\]